# Decentralized CBF-based Safety Filters for Collision Avoidance of Cooperative Missile Systems with Input Constraints

### v2

@@ I Introduction
In cooperative missile guidance, these challenges are particularly acute. Multiple interceptors must engage one or more targets, often with coordinated timing requirements to maximize success [1, 2]. At the engagement level, cooperative guidance schemes typically combine high-level target assignment with local trajectory generation. While effective for distributing interceptors, such schemes often neglect detailed interactions between neighboring effectors. As a result, even optimally coordinated assignments may yield conflicting trajectories, where interceptors converge and risk mid-air collisions despite pursuing distinct targets [3]. The central difficulty lies in bridging the gap between discrete assignment decisions at the mission-planning level and continuous-time collision avoidance during execution.
This challenge extends to safety-critical multi-agent systems more broadly, where decentralized agents must coordinate under limited communication and sensing while respecting hard safety constraints. Control Barrier Functions (CBFs) provide a principled framework for enforcing safety by certifying forward invariance of safe sets [4]. Recent work has applied barrier-function methods to real-time control, enabling constraint-satisfying modifications of nominal inputs [5]. However, most formulations assume geometric braking or idealized conditions and do not systematically address bounded inputs or feasibility issues arising in simultaneous multi-agent interactions.
A key difficulty is the “leaky corner problem,” which occurs when pairwise barrier sets intersect such that no admissible control input satisfies all constraints. Prior work has shown that the intersection of individual pairwise safe sets does not generally represent the true joint safe set, and identifying such infeasible regions without full reachability analysis remains an open problem [6]. Despite recognition of this issue, existing approaches either rely on computationally expensive global methods or neglect it entirely, limiting applicability to real-time aerospace systems.
This paper proposes a decentralized, event-triggered safety filter for cooperative guidance based on robust control barrier functions (RCBFs) [7]. The formulation explicitly incorporates input constraints in a high-relative-degree setting and ensures feasible safety decisions even with multiple conflicting agents. Each effector solves a local quadratic program (QP) that modifies its nominal guidance command only when necessary, using information from a bounded neighborhood defined by range and zero-effort miss (ZEM) criteria. To handle feasibility challenges from simultaneous constraints, we introduce a slack-variable prioritization mechanism that allocates relaxation selectively to less critical neighbors, ensuring imminent threats remain strictly enforced (see Fig. 1).
@@ II Preliminaries
The following standard definitions are recalled [8, 9]:
@@
The objective is to guarantee forward invariance of a subset of $\mathscr{S}(t)$ in order to satisfy the constraints described by $h$, despite input limitations and disturbances. This problem is studied in detail in [7], where a solution is provided through robust control barrier functions (RCBFs). The following definition is adapted from Theorem 9 in [7].
Theorem 1 ([7, Theorem 9]).
@@ IV Decentralized Robust CBF-based QP Formulation for Collision Avoidance
where $r_{\text{neigh},i}>0$ denotes the maximum interaction radius within which safety constraints are actively considered. This localized filtering reduces the number of constraints each agent must evaluate, enabling real-time feasibility in large-scale systems, without compromising safety for a sufficiently large neighborhood $\mathscr{N}_{i}$ defined via $r_{\text{neigh},i}$ [5].
@@ V Criticality-Based Prioritization under Conflicting Constraints
For the proposed decentralized, event-triggered safety filter framework, each agent autonomously evaluates and activates pairwise barrier-function constraints depending on the local interaction state. While individual pairwise constraints rigorously certify safety in isolation, their simultaneous enforcement does not automatically extend to the multi-agent setting. In particular, situations may arise where multiple constraints are active at once and form mutually restrictive conditions. A representative example is the so-called leaky-corner problem, in which the intersection of several pairwise safe sets leaves no admissible input, even though the system is physically capable of maintaining safe separation. More broadly, the lack of compositionality between local certificates and the global safe set can induce temporary deadlocks that undermine feasibility without necessarily reflecting true collision risk [6].
@@
Here, the auxiliary variables $\delta_{ij}\geq 0$ represent constraint relaxations, and the weights $w_{ij}>0$ encode interaction priorities. Large weights make violations costly, biasing the optimizer to enforce critical constraints strictly, while smaller weights allow limited relaxation on less critical ones. From a multi-objective perspective, it could be argued that the solution is Pareto optimal, in the sense that the approach seeks to avoid constraint violations whenever possible [10]. When violations are unavoidable, no single violation can be reduced without either worsening another or deviating further from the nominal guidance command. In such cases, residual violations are applied preferentially to less critical constraints. In the considered case, priorities are assigned in a state-dependent manner:
@@ VI Simulation Results
The comparison clearly shows that without the safety filter collisions can occur even when target assignments are distinct, while with the filter all effectors remain safe and complete their assigned interceptions. Furthermore, the corrective action of the safety filter is localized and minimal, intervening only when conflicts arise, and preserving the nominal performance for effectors not directly involved in potential collisions. This ensures that the proposed approach provides both safety and efficiency, making it suitable for large-scale cooperative aerospace scenarios.

### v3

@@ I Introduction
In cooperative missile guidance, these challenges are particularly acute. Multiple interceptors must engage one or more targets, often with coordinated timing requirements to maximize success [LI2024109212, Li2023]. At the engagement level, cooperative guidance schemes typically combine high-level target assignment with local trajectory generation. While effective for distributing interceptors, such schemes often neglect detailed interactions between neighboring effectors. As a result, even optimally coordinated assignments may yield conflicting trajectories, where interceptors converge and risk mid-air collisions despite pursuing distinct targets [Lv2024]. The central difficulty lies in bridging the gap between discrete assignment decisions at the mission-planning level and continuous-time collision avoidance during execution.
This challenge extends to safety-critical multi-agent systems more broadly, where decentralized agents must coordinate under limited communication and sensing while respecting hard safety constraints. Control Barrier Functions (CBFs) provide a principled framework for enforcing safety by certifying forward invariance of safe sets [ames2016control]. Recent work has applied barrier-function methods to real-time control, enabling constraint-satisfying modifications of nominal inputs [BORRMANN2015UAV]. However, most formulations assume geometric braking or idealized conditions and do not systematically address bounded inputs or feasibility issues arising in simultaneous multi-agent interactions.
A key difficulty is the “leaky corner problem,” which occurs when pairwise barrier sets intersect such that no admissible control input satisfies all constraints. Prior work has shown that the intersection of individual pairwise safe sets does not generally represent the true joint safe set, and identifying such infeasible regions without full reachability analysis remains an open problem [Aloor_2025]. Despite recognition of this issue, existing approaches either rely on computationally expensive global methods or neglect it entirely, limiting applicability to real-time aerospace systems.
This paper proposes a decentralized, event-triggered safety filter for cooperative guidance based on robust control barrier functions (RCBFs) [breeden2023robust]. The formulation explicitly incorporates input constraints in a high-relative-degree setting and ensures feasible safety decisions even with multiple conflicting agents. Each effector solves a local quadratic program (QP) that modifies its nominal guidance command only when necessary, using information from a bounded neighborhood defined by range and zero-effort miss (ZEM) criteria. To handle feasibility challenges from simultaneous constraints, we introduce a slack-variable prioritization mechanism that allocates relaxation selectively to less critical neighbors, ensuring imminent threats remain strictly enforced (see Fig. 1).
@@ II Preliminaries
The following standard definitions are recalled [Nagumo_1942, Blanchini_1999]:
@@
The objective is to guarantee forward invariance of a subset of $\mathscr{S}(t)$ in order to satisfy the constraints described by $h$, despite input limitations and disturbances. This problem is studied in detail in [breeden2023robust], where a solution is provided through robust control barrier functions (RCBFs). The following definition is adapted from Theorem 9 in [breeden2023robust].
Theorem 1 ([breeden2023robust, Theorem 9]).
@@ IV Decentralized Robust CBF-based QP Formulation for Collision Avoidance
where $r_{\text{neigh},i}>0$ denotes the maximum interaction radius within which safety constraints are actively considered. This localized filtering reduces the number of constraints each agent must evaluate, enabling real-time feasibility in large-scale systems, without compromising safety for a sufficiently large neighborhood $\mathscr{N}_{i}$ defined via $r_{\text{neigh},i}$ [BORRMANN2015UAV].
@@ V Criticality-Based Prioritization under Conflicting Constraints
For the proposed decentralized, event-triggered safety filter framework, each agent autonomously evaluates and activates pairwise barrier-function constraints depending on the local interaction state. While individual pairwise constraints rigorously certify safety in isolation, their simultaneous enforcement does not automatically extend to the multi-agent setting. In particular, situations may arise where multiple constraints are active at once and form mutually restrictive conditions. A representative example is the so-called leaky-corner problem, in which the intersection of several pairwise safe sets leaves no admissible input, even though the system is physically capable of maintaining safe separation. More broadly, the lack of compositionality between local certificates and the global safe set can induce temporary deadlocks that undermine feasibility without necessarily reflecting true collision risk [Aloor_2025].
@@
Here, the auxiliary variables $\delta_{ij}\geq 0$ represent constraint relaxations, and the weights $w_{ij}>0$ encode interaction priorities. Large weights make violations costly, biasing the optimizer to enforce critical constraints strictly, while smaller weights allow limited relaxation on less critical ones. From a multi-objective perspective, it could be argued that the solution is Pareto optimal, in the sense that the approach seeks to avoid constraint violations whenever possible [Cao_2025]. When violations are unavoidable, no single violation can be reduced without either worsening another or deviating further from the nominal guidance command. In such cases, residual violations are applied preferentially to less critical constraints. In the considered case, priorities are assigned in a state-dependent manner:
@@ VI Simulation Results
The comparison clearly shows that without the safety filter collisions can occur, even when target assignments are distinct, whereas with the filter all effectors remain safe and complete their assigned interceptions. Furthermore, the corrective action of the safety filter is localized and minimal, intervening only when conflicts arise, and preserving the nominal performance for effectors not directly involved in potential collisions. This ensures that the proposed approach provides both safety and efficiency, making it suitable for large-scale cooperative aerospace scenarios.